OpenStudy (anonymous):

so i have a question, how do i write the standard form of this equation of a line that is paralell to the given line? 2x-y+6=0 , (0,4)

OpenStudy (anonymous):

the given line can be written y = 2x + 6 which is the slope / intercept form the slope is coefficient of x = 2 the line parallel must also have slope 2 it must also pass thru point (0,4). using the standard form y-y1 = m(x-x1) where m is slope and x1 = 0 y1 = 4 y-4 = 2(x-0) y - 4 = 2x y - 2x - 4 = 0

OpenStudy (amistre64):

remove the constant and calibrate with the given point for a new constant; the rest of it stays the same

OpenStudy (amistre64):

while jimmys is fine; some frown upon not haveing the x first and it being negative if it is

OpenStudy (amistre64):

2x -y +4 = 0

OpenStudy (anonymous):

thank you all for all he help i might be posting some more on here before the night is over